OpenStudy (mayaal):

Please Help! will give out medals! Let f(x)=x^4+2x^2, find the equation of the line tangent f where f'(x)=1

OpenStudy (peachpi):

The slope of the tangent is 1. Find f'(x) and then set it equal to 1 to find the x-coordinate. Then you should be able to solve for the y and write the equation of the line

OpenStudy (mayaal):

I can't solve for x when i set f'(x) equal to 1

OpenStudy (peachpi):

what did you get for f'(x)?

OpenStudy (mayaal):

4x^3+4x

OpenStudy (peachpi):

There's not a nice way to solve it algebraically. You have to try it with a numerical method or a calculator

OpenStudy (peachpi):

try Newton's method if you can't use a calculator

OpenStudy (sooobored):

er, what? f'(x) is the slope of the line assuming you want to find the line tangent at x=1 then you can determine the coordinate point from f(1)= ... f'(1) will give slope, assuming you wnat a linear line, use point slope formula

OpenStudy (mayaal):

but what are the coordinates for the point slope form? @sooobored

OpenStudy (mayaal):

@mathmate

OpenStudy (sooobored):

f(x)=x^4+2x^2 f(1)=1^4+2*1^2 f(1)=1+2 f(1)=3 (1,3)

OpenStudy (mayaal):

the answer is y=x-0.122 but i want an explanation as to how to get this answer

OpenStudy (sooobored):

oh, i think i misread the problem

OpenStudy (sooobored):

find when the tangent line when f'(x)=1 we know f'(x)= 4x^3+4x so we want to find x when 1=4x^3+4x

OpenStudy (mayaal):

yeah

OpenStudy (sooobored):

4x^3+4x-1=0 so find the cubic roots of x

OpenStudy (mayaal):

That is where i am having trouble, I am unable to find x

OpenStudy (sooobored):

fyi, its not a very nice problem, the root turns out to be x~0.23673, other two roots are imaginary

OpenStudy (mayaal):

so then we can plug in x in f(x) to find the y value, right?

OpenStudy (sooobored):

yea

OpenStudy (sooobored):

cubic formula http://www.math.vanderbilt.edu/~schectex/courses/cubic/

OpenStudy (mayaal):

Thankyou so much, I finally got the answer :D